# Supplemetary Material for the paper

### v1

@@ Appendix C Supporting Lemmas and the Proof of the Main Theorem
To quantify the proximity between the estimated parameters $(\mathbf{g}(t),\mathbf{V}(t))$ and reference parameters $(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))$, we use the squared distance defined in (82), namely $\Big\|(\mathbf{g}(t),\mathbf{V}(t))\left.-(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))\right\|^{2}$.
Bounding the Initial Distance
We first analyze the initial distance between the parameters for any row $i$ that is outside the support set $(i\notin\mathcal{S}_{K})$. For these rows, we recall that the estimated parameters are initialized as $g_{i}(0)=\alpha_{g}$ and $V_{il}(0)=\alpha_{V}$, while the reference parameters are $\tilde{g}_{i}(0)=0$ and $\tilde{V}_{il}(0)=0$ $\forall l\in[L]$.
The distance contributed by the non-support points ($i\notin\mathcal{S}_{K}$) is:
$\sum\limits_{l\in[L]}\Big($ $\left|g_{i}(0)-\tilde{g}_{i}(0)\right|^{2}+\left|V_{il}(0)-\tilde{V}_{il}(0)\right|^{2}\Big)$ (200)
$=L\alpha_{g}^{2}+L\alpha_{V}^{2}.$ (201)
Using the balancedness condition at $t=0$ $\left(\frac{1}{2}g_{i}^{2}(t)=\sum\limits_{j}V_{ij}^{2}(t)\right))$ and $\alpha_{g}^{2}=2L\alpha_{V}^{2}$, this simplifies to
$\sum\limits_{l\in[L]}\Big($ $\left|g_{i}(0)-\tilde{g}_{i}(0)\right|^{2}+\left|V_{il}(0)-\tilde{V}_{il}(0)\right|^{2}\Big)$
$=\left({2L}+1\right)L\alpha_{V}^{2}.$ (202)
To control how this distance evolves over time, we use razin2021implicit. This lemma requires us to set the initial distance to be bounded by a term related to the final approximation error $\epsilon$ and the smoothness constant $\beta$, which we ensure by imposing a condition on the initialization scales, which is:
$\sum\limits_{l\in[L]}\left(\left|g_{i}(0)-\tilde{g}_{i}(0)\right|^{2}+\left|V_{il}(0)-\tilde{V}_{il}(0)\right|^{2}\right)\leq\frac{1}{N}\text{exp}(-2\beta T)$ (203)
Then we have $\left({2L}+1\right)L\alpha_{V}^{2}\leq\frac{1}{N}\text{exp}(-2\beta T)$, rearranging this inequality gives us the upper bound on the initialization as:
$\alpha_{V}\leq\frac{\text{exp}(-\beta T)}{\sqrt{LN\left({2L}+1\right)}}.$ (204)
The distance contributed by the support set $(i\in\mathcal{S}_{K})$ at initialization $t=0$ is:
$\sum\limits_{l\in[L]}\left(\left|g_{i}(0)-\tilde{g}_{i}(0)\right|^{2}+\left|V_{il}(0)-\tilde{V}_{il}(0)\right|^{2}\right)$ $=L(\alpha_{g}^{2}+\rho^{2/3}-2\alpha_{g}\rho^{1/3}+\alpha_{V}^{2}+\rho^{2/3}-2\alpha_{V}\rho^{1/3})$ (205)
$=L((2L+1)\alpha_{V}^{2}+2\rho^{2/3}-2(\sqrt{2L}+1)\alpha_{V}\rho^{1/3}).$ (206)
Let $0<\rho\leq\alpha_{V}^{3}$. By applying the bound from (203) we have:
$\left(L+\frac{1}{2}\right)\alpha_{V}^{2}+\rho^{2/3}-\left(\sqrt{2L}+1\right)\alpha_{V}\rho^{1/3}$ $\leq\frac{1}{2LN}\text{exp}\left(-2\beta T\right)$ (207)
$\rho^{2/3}-\left(\sqrt{2L}+1\right)\alpha_{V}\rho^{1/3}+\left(\sqrt{\frac{L}{2}}+\frac{1}{2}\right)^{2}\alpha_{V}^{2}$ $\leq\frac{1}{2LN}\text{exp}\left(-2\beta T\right)-\left(L+\frac{1}{2}\right)\alpha_{V}^{2}+\left(\sqrt{\frac{L}{2}}+\frac{1}{2}\right)^{2}\alpha_{V}^{2}$ (208)
$\left(\rho^{1/3}-\left(\sqrt{\frac{L}{2}}+\frac{1}{2}\right)\alpha_{V}\right)^{2}$ $\leq\frac{1}{2LN}\text{exp}\left(-2\beta T\right)-\frac{1}{4}\alpha_{V}^{2}\left(\sqrt{2L}-1\right)^{2}.$ (209)
For a real solution for $\rho^{1/3}$ to exist, the right-hand side of this inequality must be non-negative. Then, let $\alpha_{V}$ be designed such that $\frac{1}{2LN}\text{exp}\left(-2\beta T\right)-\frac{1}{4}\alpha_{V}^{2}\left(\sqrt{2L}-1\right)^{2}\geq 0$ holds. Thus,
$\alpha_{V}\leq\frac{\sqrt{2}}{\sqrt{LN}({\sqrt{2L}-1})}\text{exp}(-\beta T).$ (210)
To ensure that the initialization conditions for rows both inside and outside the support set are met, $\alpha_{V}$ must satisfy the stricter of the two derived upper bounds ((204) and (204)). This leads to the initialization condition:
$\alpha_{V}\leq\frac{1}{\sqrt{LN}}\text{exp}(-\beta T)\text{min}\left\{\frac{\sqrt{2}}{\sqrt{2L}-1},\frac{1}{\sqrt{\left({2L}+1\right)}}\right\}.$ (211)
Since $\frac{\sqrt{2}}{\sqrt{2L}-1}>\frac{1}{\sqrt{\left({2L}+1\right)}}$ for all $L\geq 1$. Note that this is still Equation (204).
Equation (209) provides a valid interval for $\rho^{1/3}$:
$\rho^{1/3}\in\left[\left(\sqrt{\frac{L}{2}}+\frac{1}{2}\right)\alpha_{V}-\epsilon_{\alpha},\left(\sqrt{\frac{L}{2}}+\frac{1}{2}\right)\alpha_{V}+\epsilon_{\alpha}\right],$ (212)
where $\epsilon_{\alpha}=\left(\frac{1}{2LN}\text{exp}\left(-2\beta T\right)-\frac{1}{4}\alpha_{V}^{2}\left(\sqrt{2L}-1\right)^{2}\right)^{1/2}$. But we have $\left(\sqrt{\frac{L}{2}}+\frac{1}{2}\right)\alpha_{V}+\epsilon_{\alpha}\geq\alpha_{V}$ in order to keep $\epsilon_{\alpha}\geq 0$. Hence the set becomes:
$\rho^{1/3}\in\left[\left(\sqrt{\frac{L}{2}}+\frac{1}{2}\right)\alpha_{V}-\epsilon_{\alpha},\alpha_{V}\right].$ (213)
For this set to be non-empty, we need lower bound to be less than or equal to the upper bound. This gives us:
$\left(\sqrt{\frac{L}{2}}+\frac{1}{2}\right)\alpha_{V}-\epsilon_{\alpha}$ $\leq\alpha_{V}.$ (214)
Rearranging, we have:
$\left(\sqrt{\frac{L}{2}}-\frac{1}{2}\right)\alpha_{V}$ $\leq\epsilon_{\alpha}.$ (215)
Rearranging both sides and substituting for $\epsilon_{\alpha}$:
$\left(\sqrt{\frac{L}{2}}-\frac{1}{2}\right)^{2}\alpha_{V}^{2}$ $\leq\frac{1}{2LN}\text{exp}\left(-2\beta T\right)-\frac{1}{4}\alpha_{V}^{2}\left(\sqrt{2L}-1\right)^{2}.$ (216)
Reorganizing,
$\alpha_{V}^{2}\left(\sqrt{2L}-1\right)^{2}$ $\leq\frac{1}{LN}\text{exp}\left(-2\beta T\right),$ (218)
$\alpha_{V}$ $\leq\frac{1}{\sqrt{LN}\left(\sqrt{2L}-1\right)}\text{exp}\left(-\beta T\right).$ (219)
This confirms that the interval for $\rho$ (equation (212) is consistent with Equation (204).
With the initialization conditions for $\alpha_{V}$ and $\rho$ established, we can bound the total initial squared distance between the parameters. A time $t=0$:
$\Big\|(\mathbf{g}(0),\mathbf{V}(0))\left.-(\tilde{\mathbf{g}}(0),\tilde{\mathbf{V}}(0))\right\|^{2}$ $=\sum_{i\in[N]}\sum\limits_{l\in[L]}\left(\left|g_{i}(0)-\tilde{g}_{i}(0)\right|^{2}+\left|V_{il}(0)-\tilde{V}_{il}(0)\right|^{2}\right),$ (220)
Splitting the contributions of support and out-side support parameters,
$\Big\|(\mathbf{g}(0),\mathbf{V}(0))\left.-(\tilde{\mathbf{g}}(0),\tilde{\mathbf{V}}(0))\right\|^{2}$ $=\sum_{i\in\mathcal{S}_{K}}\sum\limits_{l\in[L]}\left(\left|g_{i}(0)-\tilde{g}_{i}(0)\right|^{2}+\left|V_{il}(0)-\tilde{V}_{il}(0)\right|^{2}\right)$
$\quad+\sum_{i\notin\mathcal{S}_{K}}\sum\limits_{l\in[L]}\left(\left|g_{i}(0)-\tilde{g}_{i}(0)\right|^{2}+\left|V_{il}(0)-\tilde{V}_{il}(0)\right|^{2}\right),$ (221)
Substituting back the equations (206) and (202),
$\Big\|(\mathbf{g}(0),\mathbf{V}(0))\left.-(\tilde{\mathbf{g}}(0),\tilde{\mathbf{V}}(0))\right\|^{2}$ $\leq KL((2L+1)\alpha_{V}^{2}+2\rho^{2/3}-2(\sqrt{2L}+1)\alpha_{V}\rho^{1/3})+(N-K)(\left({2L}+1\right)L\alpha_{V}^{2}),$ (222)
$\leq KL(2L+1)\alpha_{V}^{2}+2KL\rho^{2/3}-2KL(\sqrt{2L}+1)\alpha_{V}\rho^{1/3}+(N-K)(\left({2L}+1\right)L\alpha_{V}^{2}),$ (223)
$\leq NL(2L+1)\alpha_{V}^{2}+2KL\rho^{2/3}-2KL(\sqrt{2L}+1)\alpha_{V}\rho^{1/3},$ (224)
to find a simpler upper bound, we can drop the negative term.
$\Big\|(\mathbf{g}(0),\mathbf{V}(0))\left.-(\tilde{\mathbf{g}}(0),\tilde{\mathbf{V}}(0))\right\|^{2}$ $\leq NL(2L+1)\alpha_{V}^{2}+2KL\rho^{2/3}.$ (225)
We have $0<\rho\leq\alpha_{V}^{3}$, which implies $\rho^{2/3}\leq\alpha_{V}^{2}$. Hence,
$\Big\|(\mathbf{g}(0),\mathbf{V}(0))\left.-(\tilde{\mathbf{g}}(0),\tilde{\mathbf{V}}(0))\right\|^{2}$ $\leq NL(2L+1)\alpha_{V}^{2}+2K\alpha_{V}^{2}$ (226)
$\leq(N(2L+1)+2K)L\alpha_{V}^{2}.$ (227)
Note that the substituting the upper bound on the value $\alpha_{V}$ from equation (204) yields:
$\Big\|(\mathbf{g}(0),\mathbf{V}(0))\left.-(\tilde{\mathbf{g}}(0),\tilde{\mathbf{V}}(0))\right\|^{2}\leq\text{exp}\left(-2\beta T\right).$
Bounding the Evolved Distance
[…]

### v2

@@ 2. Case II:
To quantify the proximity between the estimated parameters $(\mathbf{g}(t),\mathbf{V}(t))$ and reference parameters $(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))$, we use the squared distance, $\Big\|(\mathbf{g}(t),\mathbf{V}(t))\left.-(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))\right\|^{2}$.
We formally define the Lyapunov function $\mathcal{E}(t)$ as:
$\mathcal{E}(t)$ $=\Big\|(\mathbf{g}(t),\mathbf{V}(t))\left.-(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))\right\|^{2}=\sum\limits_{i\in[N]}\sum\limits_{l\in[L]}\Big(\big|g_{i}(t)-\tilde{g}_{i}(t)\big|^{2}+\big|V_{il}(t)-\tilde{V}_{il}(t)\big|^{2}\Big)$ (153)
We decompose this into summation of independent row-wise components, $\mathcal{E}_{i}(t)$, corresponding to each row index $i\in[N]$:
$\mathcal{E}(t)$ $=\sum\limits_{i\in[N]}\mathcal{E}_{i}(t)$ (154)
$\text{where }\mathcal{E}_{i}(t)$ $=\sum\limits_{l\in[L]}\Big(\big|g_{i}(t)-\tilde{g}_{i}(t)\big|^{2}+\big|V_{il}(t)-\tilde{V}_{il}(t)\big|^{2}\Big),\quad\forall i\in[N]$ (155)
C.1.1 Bounding the error on non-support indices $(i\notin\mathcal{S}_{K})$
We begin by analyzing the error growth on the non-support indices/rows $\big(i\notin\mathcal{S}_{K}\big)$. For these rows, the reference trajectory is zero (by construction $\tilde{g}_{i}(0)=0,\tilde{V}_{il}(0)=0$ and its subsequent updates $\tilde{g}_{i}(t)=0,\tilde{V}_{il}(t)=0$ as a consequence of Lemma C.1). Consequently, the row-wise error term simplifies to the energy of estimated trajectory alone.
$\mathcal{E}_{i}(t)=\mathcal{E}_{e_{i}}(t)$ $=\sum\limits_{l\in[L]}\Big(g_{i}(t)^{2}+V_{il}(t)^{2}\Big)$ (156)
From perfect row-balancedness property, i.e., $\Big(\frac{1}{2}g_{i}^{2}(t)=\sum_{l\in[L]}V_{il}^{2}(t)\Big)$, error in terms of $|g_{i}(t)|$ is expressed as
$\mathcal{E}_{e_{i}}(t)$ $=\Bigg(L+\frac{1}{2}\Bigg)|g_{i}(t)|^{2}$ (157)
At initialization $(t=0)$, the error is given by:
$\mathcal{E}_{{e}_{i}}(0)$ $=\Bigg(L+\frac{1}{2}\Bigg)\alpha_{g}^{2}=2L\Bigg(L+\frac{1}{2}\Bigg)\alpha_{V}^{2}$ (158)
Next, to bound the growth of error, we differentiate with respect to time on both sides:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{e_{i}}(t)$ $=\sum\limits_{l\in[L]}\bigg(\frac{\mathrm{d}}{\mathrm{d}t}g_{i}(t)^{2}+\frac{\mathrm{d}}{\mathrm{d}t}V_{il}(t)^{2}\bigg)$ (159)
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{e_{i}}(t)$ $=\bigg(L\frac{\mathrm{d}}{\mathrm{d}t}g_{i}(t)^{2}+\frac{\mathrm{d}}{\mathrm{d}t}\sum\limits_{l\in[L]}V_{il}(t)^{2}\bigg)$ (160)
From equation (42), (i.e. $\Big(\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}g_{i}^{2}(t)=\frac{\mathrm{d}}{\mathrm{d}t}\|\mathbf{V}_{i:}(t)\|_{2}^{2}=\frac{\mathrm{d}}{\mathrm{d}t}\sum\limits_{l\in[L]}V_{il}^{2}\Big)$
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{e_{i}}(t)$ $=\Bigg(L+\frac{1}{2}\Bigg)\bigg(\frac{\mathrm{d}}{\mathrm{d}t}g_{i}(t)^{2}\bigg)$ (161)
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{e_{i}}(t)$ $=2\Bigg(L+\frac{1}{2}\Bigg)\bigg(g_{i}(t)\frac{\mathrm{d}}{\mathrm{d}t}g_{i}(t)\bigg)$ (162)
Substituting the gradient flow dynamics for $g_{i}(t)$ from equation (34) $\Big(\frac{\mathrm{d}}{\mathrm{d}t}{g}_{i}(t)=4g_{i}(t)\sum\limits_{j\in[L]}\Lambda_{ij}(t)V_{lj}(t)\Big)$:
$=\bigg(L+\frac{1}{2}\bigg)8g_{i}^{2}(t)\sum\limits_{l\in[L]}\Lambda_{il}(t)V_{il}(t)$ (163)
Applying Cauchy-Schwarz inequality yields the following upper-bound:
$\leq\Bigg(L+\frac{1}{2}\Bigg)8g_{i}^{2}(t)\big\|\mathbf{\Lambda}_{i:}(t)\big\|_{2}\|\mathbf{V}_{i:}(t)\big\|_{2}$ (164)
To establish an upper bound on the residual term $\left\|\mathbf{\Lambda}_{i:}(t)\right\|_{2}$, we leverage the monotonic decrease of the loss function (Ji and Telgarsky, 2019).
$\|\mathbf{\Lambda}_{i:}(t)\|_{2}$ $=\|\mathbf{A}^{\top}_{i:}\left(\mathbf{Y}-\mathbf{A}\mathbf{X}(t)\right)\|_{2}\overset{(a)}{\leq}\|\mathbf{A}^{\top}_{i:}\|_{2}\|\left(\mathbf{Y}-\mathbf{A}\mathbf{X}(t)\right)\|_{F}\overset{(b)}{\leq}\|\left(\mathbf{Y}-\mathbf{A}\mathbf{X}(0)\right)\|_{F}$ (165)
${\leq}\|\mathbf{Y}\|_{F}+\|\mathbf{A}\|_{F}\|\mathbf{X}(0)\|_{F}\overset{(c)}{\leq}\|\mathbf{Y}\|_{F}+N\sqrt{L}\alpha_{g}^{2}\alpha_{V}=\|\mathbf{Y}\|_{F}+2L^{3/2}N\alpha_{V}^{3}:=C_{\Lambda}$ (166)
Where $\overset{(a)}{}$ follows from sub-multiplicativity (Belitskii and others, 2013) of the Frobenius norm (which is a direct consequence of the Cauchy-Schwarz inequality applied column-wise), $\overset{(b)}{}$ from the $\ell_{2}$ normalization of columns of $\mathbf{A}$ and the monotonic decrease of loss and $\overset{(c)}{}$ from equation (113) Recall that the gradient flow preserves row-wise balancedness, $\Big(\frac{1}{2}g_{i}(t)^{2}=\sum\limits_{j\in[L]}V_{ij}(t)^{2}\Big)$. Substituting the above derived bound for $\|\mathbf{\Lambda}_{i:}(t)\|_{2}$ and balancedness into equation (164), we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{e_{i}}(t)$ $\leq 4\sqrt{2}\Bigg(L+\frac{1}{2}\Bigg)C_{\Lambda}|g_{i}(t)|^{3}$ (167)
From equation (158), we can substitute for $|g_{i}(t)|=\big(L+\frac{1}{2}\big)^{-1/2}\mathcal{E}_{e_{i}}(t)^{1/2}$. Then the bound becomes
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{e_{i}}(t)$ $\leq 4\sqrt{2}\Bigg(L+\frac{1}{2}\Bigg)^{-\frac{1}{2}}C_{\Lambda}\mathcal{E}_{e_{i}}^{3/2}(t)=C_{\Lambda}^{\prime}\mathcal{E}_{e_{i}}^{3/2}(t)$ (168)
where, $C_{\Lambda}^{\prime}:=4\sqrt{2}\Bigg(L+\frac{1}{2}\Bigg)^{-\frac{1}{2}}C_{\Lambda}$. Rearranging the terms to separate the variables $\mathcal{E}_{e_{i}}$ and $t$ in equation (168), we obtain:
$\mathcal{E}_{e_{i}}^{-3/2}(t)\frac{\mathrm{d}\mathcal{E}_{e_{i}}(t)}{\mathrm{d}t}$ $\leq C_{\Lambda}^{\prime}$ (169)
$\int_{0}^{t}\mathcal{E}_{e_{i}}^{-3/2}(\tau)\frac{\mathrm{d}\mathcal{E}_{e_{i}}(\tau)}{\mathrm{d}\tau}\mathrm{d}\tau$ $\leq\int_{0}^{t}C_{\Lambda}^{\prime}\mathrm{d}\tau$
$\int_{\mathcal{E}_{e_{i}}(0)}^{\mathcal{E}_{e_{i}}(t)}u^{-3/2}\mathrm{d}u$ $\leq C_{\Lambda}^{\prime}t$ (170)
$\left[-2u^{-1/2}\right]_{\mathcal{E}_{e_{i}}(0)}^{\mathcal{E}_{e_{i}}(t)}$ $\leq C_{\Lambda}^{\prime}t$
$-2\mathcal{E}_{e_{i}}^{-1/2}(t)+2\mathcal{E}_{e_{i}}^{-1/2}(0)$ $\leq C_{\Lambda}^{\prime}t$ (171)
Isolating the term involving $\mathcal{E}_{e_{i}}(t)$:
$\mathcal{E}_{e_{i}}^{-1/2}(t)$ $\geq\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t$ (172)
Taking the reciprocal and squaring both sides yields the upper bound:
$\mathcal{E}_{e_{i}}(t)$ $\leq\frac{1}{\Big(\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t\Big)^{2}}$ (173)
C.1.2 Bounding the error on support indices $(i\in\mathcal{S}_{K})$
We now extend our analysis to the rows in the support set $\mathcal{S}_{K}$. Unlike the non-support indices where the reference trajectory remains zero, here the reference trajectory $\tilde{\mathbf{X}}(t)$ evolves non-trivially according to the gradient flow dynamics. To bound the divergence for these indices, we decompose the total error $\mathcal{E}_{i}(t)$ into three distinct components: the energy of the estimated trajectory $\mathcal{E}_{e_{i}}(t)$, the energy of the reference trajectory $\mathcal{E}_{r_{i}}(t)$, and a cross-term $\mathcal{E}_{x_{i}}(t)$. For each row, $i\in\mathcal{S}_{k}$:
$\mathcal{E}_{i}(t)$ $=\sum\limits_{l\in[L]}\Big(\big(g_{i}(t)-\tilde{g}_{i}(t)\big)^{2}+\big(V_{il}(t)-\tilde{V}_{il}(t)\big)^{2}\Big)$ (174)
Expanding the squared terms and regrouping the components, we obtain:
$=\sum\limits_{l\in[L]}\Big(g_{i}^{2}(t)+\tilde{g}_{i}^{2}(t)-2g_{i}(t)\tilde{g}_{i}(t)+V_{il}^{2}(t)+\tilde{V}_{il}^{2}(t)-2V_{il}(t)\tilde{V}_{il}(t)\Big)$ (175)
$=\mathcal{E}_{e_{i}}(t)+\mathcal{E}_{r_{i}}(t)-2\mathcal{E}_{x_{i}}(t)$ (176)
where the reference energy $\mathcal{E}_{r_{i}}(t)$ and the cross-term energy $\mathcal{E}_{x_{i}}(t)$ are defined as:
$\mathcal{E}_{r_{i}}(t)$ $=\sum\limits_{l\in[L]}\Big(\tilde{g}_{i}(t)^{2}+\tilde{V}_{il}(t)^{2}\Big),$ (177)
$\mathcal{E}_{x_{i}}(t)$ $=\sum\limits_{l\in[L]}\Big(g_{i}(t)\tilde{g}_{i}(t)+V_{il}(t)\tilde{V}_{il}(t)\Big).$ (178)
First we bound the energy of the reference trajectory, $\mathcal{E}_{r_{i}}(t)$.
Similar to the approach for the estimated trajectory, we utilize the perfect row-balancedness to express the energy in terms of the scaling factor $\tilde{g}_{i}(t)$:
$\mathcal{E}_{r_{i}}(t)$ $=\sum\limits_{l\in[L]}\Big(\tilde{g}_{i}(t)^{2}+\tilde{V}_{il}(t)^{2}\Big)$ (179)
$\mathcal{E}_{r_{i}}(t)$ $=\Bigg(L+\frac{1}{2}\Bigg)\tilde{g}_{i}(t)^{2}$ (180)
At time $t=0$, the error is
$\mathcal{E}_{r_{i}}(0)$ $=\left(2L+1\right)\rho^{2/3}$ (181)
To bound the growth of energy, we differentiate with respect to time. Substituting the gradient flow dynamics (Lemma C.1) and applying the Cauchy-Schwarz inequality yields the following upper bound:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{r_{i}}(t)$ $=\Bigg(L+\frac{1}{2}\Bigg)\bigg(\frac{\mathrm{d}}{\mathrm{d}t}\tilde{g}_{i}(t)^{2}\bigg)$ (182)
$\leq\Bigg(L+\frac{1}{2}\Bigg)8\tilde{g}_{i}^{2}(t)\big\|\mathbf{\tilde{\Lambda}}_{i:}(t)\big\|_{2}\|\tilde{\mathbf{V}}_{i:}(t)\big\|_{2}$ (183)
where we have utilized $\Big(\frac{\mathrm{d}}{\mathrm{d}t}\tilde{g}_{i}(t)=4\tilde{g}_{i}(t)\sum\limits_{j\in[L]}\tilde{\Lambda}_{ij}(t)\tilde{V}_{lj}(t)\Big)$.
Now, we establish an upper bound on the residual term $\|\tilde{\mathbf{\Lambda}}_{i:}(t)\|_{2}$. Proceeding similarly to the derivation of $C_{\Lambda}$, we leverage the monotonic decrease of the loss function to bound the residual by its initialization values:
$\|\tilde{\mathbf{\Lambda}}_{i:}(t)\|_{2}$ $\leq\|\mathbf{Y}\|_{F}+\|\mathbf{A}\|_{F}\|\tilde{\mathbf{X}}(0)\|_{F}\leq\|\mathbf{Y}\|_{F}+\sqrt{K}N\rho:=\tilde{C}_{\Lambda}$ (184)
Substituting the above bound and perfect row-balancedness, we arrive at the differential inequality:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{r_{i}}(t)$ $\leq 4\sqrt{2}\Bigg(L+\frac{1}{2}\Bigg)C_{\Lambda}|\tilde{g}_{i}(t)|^{3}$ (185)
Using equation (180) we can substitute $|\tilde{g}_{i}(t)|=\big(L+\frac{1}{2}\big)^{-1/2}\mathcal{E}_{r_{i}}(t)^{1/2}$.
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{r_{i}}(t)$ $\leq 4\sqrt{2}\Bigg(L+\frac{1}{2}\Bigg)^{-\frac{1}{2}}\tilde{C}_{\Lambda}\mathcal{E}_{r_{i}}^{3/2}(t)=\tilde{C}_{\Lambda}^{\prime}\mathcal{E}_{r_{i}}^{3/2}(t)$ (186)
where, $\tilde{C}_{\Lambda}^{\prime}:=4\sqrt{2}\Bigg(L+\frac{1}{2}\Bigg)^{-\frac{1}{2}}\tilde{C}_{\Lambda}$
The above differential inequality is structurally identical to the one for the estimated trajectory solved in the sub-section C.1.1. Integrating it over the interval $[0,t]$ immediately yields the analogous upper bound:
$\mathcal{E}_{r_{i}}(t)\leq\frac{1}{\Big(\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t\Big)^{2}}.$ (187)
Bounding the Cross-Term $\left(\mathcal{E}_{x_{i}}\right)$:
Having established upper bounds for the energy of both the estimated and reference trajectories, we now bound the cross-term $\mathcal{E}_{x_{i}}(t)$, which represents the alignment between the estimated and reference trajectories.
Recall the decomposition of the total error $\mathcal{E}_{i}(t)$, which we restate here for brevity:
$\mathcal{E}_{i}(t)=\mathcal{E}_{e_{i}}(t)+\mathcal{E}_{r_{i}}(t)-2\mathcal{E}_{x_{i}}(t).$ (176)
From this relation, to establish an upper bound on the total error $\mathcal{E}_{i}(t)$, we require a lower bound on the cross-term $\mathcal{E}_{x_{i}}(t)$.
$\mathcal{E}_{x_{i}}(t)$ $=\sum\limits_{l\in[L]}\left(g_{i}(t)\tilde{g}_{i}(t)+V_{il}(t)\tilde{V}_{il}(t)\right)$ (188)
Recall that at $t=0$, the initializations are $V_{ij}(0)=\alpha_{V},g_{i}(0)=\alpha_{g}=\sqrt{2L}\alpha_{V},\tilde{g}_{i}(0)=\sqrt{2}\rho^{1/3}\text{ and }\tilde{V}_{ij}(0)=\frac{\rho^{1/3}}{\sqrt{L}}$. The initial cross-term energy is
$\mathcal{E}_{x_{i}}(0)$ $=\sqrt{L}(2L+1)\alpha_{V}\rho^{1/3}$ (189)
Differentiating with respect to time and substituting the gradient flow dynamics as before:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{x_{i}}(t)$ $=\sum\limits_{l\in[L]}g_{i}(t)\left(\frac{\mathrm{d}}{\mathrm{d}t}\tilde{g}_{i}(t)\right)+\left(\frac{\mathrm{d}}{\mathrm{d}t}g_{i}(t)\right)\tilde{g}_{i}(t)+V_{il}(t)\left(\frac{\mathrm{d}}{\mathrm{d}t}\tilde{V}_{il}(t)\right)+\left(\frac{\mathrm{d}}{\mathrm{d}t}V_{il}(t)\right)\tilde{V}_{il}(t)$ (190)
$=\sum\limits_{l\in[L]}\left(4g_{i}(t)\tilde{g}_{i}(t)\sum\limits_{j\in[L]}\tilde{\Lambda}_{ij}(t)\tilde{V}_{ij}(t)+4g_{i}(t)\tilde{g}_{i}(t)\sum\limits_{j\in[L]}\Lambda_{ij}(t)V_{ij}(t)+2\tilde{g}_{i}^{2}(t)V_{il}(t)\tilde{\Lambda}_{il}(t)+2g_{i}^{2}(t)\Lambda_{il}(t)\tilde{V}_{il}(t)\right)$ (191)
Summing over $l$ allows us to group terms as
$=4Lg_{i}(t)\tilde{g}_{i}(t)\left(\sum\limits_{j\in[L]}\tilde{\Lambda}_{ij}(t)\tilde{V}_{ij}(t)+\sum\limits_{j\in[L]}\Lambda_{ij}(t)V_{ij}(t)\right)+2|\tilde{g}_{i}(t)|^{2}\sum\limits_{l\in[L]}V_{il}(t)\tilde{\Lambda}_{il}(t)+2|g_{i}(t)|^{2}\sum\limits_{l\in[L]}\Lambda_{il}(t)\tilde{V}_{il}(t)$ (192)
Rewriting in terms of inner products:
$=4Lg_{i}(t)\tilde{g}_{i}(t)\left(\left<\tilde{\mathbf{\Lambda}}_{i:}(t)\tilde{\mathbf{V}}_{i:}(t)\right>+\Big<\mathbf{\Lambda}_{i:}(t)\mathbf{V}_{i:}(t)\Big>\right)+2\tilde{g}_{i}^{2}(t)\left<\mathbf{V}_{i:}(t)\tilde{\Lambda}_{i:}(t)\right>+2g_{i}^{2}(t)\left<\Lambda_{i:}(t)\tilde{V}_{i:}(t)\right$> (193)
Applying the Cauchy-Schwarz-Bunjakowski inequality (Teschl, 2012) $\left(\left|\langle A,B\rangle\right|\leq\|A\|\|B\|\implies-\|A\|\|B\|\leq\langle A,B\rangle\right)$ :
$\geq-4L|g_{i}(t)||\tilde{g}_{i}(t)|\left(\|\tilde{\mathbf{\Lambda}}_{i:}(t)\|\|\tilde{\mathbf{V}}_{i:}(t)\|+\|\mathbf{\Lambda}_{i:}(t)\|\|\mathbf{V}_{i:}(t)\|\right)-2\tilde{g}_{i}^{2}(t)\|\mathbf{V}_{i:}(t)\|\|\tilde{\Lambda}_{i:}(t)\|-2g_{i}^{2}(t)\|\Lambda_{i:}(t)\|\|\tilde{V}_{i:}(t)\|$ (194)
Substituting the bounds $\|\mathbf{\Lambda}_{i:}\|\leq C_{\Lambda}$ and $\|\tilde{\mathbf{\Lambda}}_{i:}\|\leq\tilde{C}_{\Lambda}$
$\geq-4L|g_{i}(t)||\tilde{g}_{i}(t)|\left(\tilde{C}_{\Lambda}(t)\|\tilde{\mathbf{V}}_{i:}(t)\|+C_{\Lambda}\|\mathbf{V}_{i:}(t)\|\right)-2\tilde{C}_{\Lambda}\tilde{g}_{i}^{2}(t)\|\mathbf{V}_{i:}(t)\|-2C_{\Lambda}(t)g_{i}^{2}(t)\|\tilde{V}_{i:}(t)\|$ (195)
Applying row-balancedness (i.e., $\sqrt{2}\|\mathbf{V}_{i:}\|=|g_{i}|$)
[…]